OpenStudy (anonymous):

Alberto measured the length of a piece of wire to be 105cm. The actual length of the wire was 120 cm. What was the percent error of Alberto's measurement?

OpenStudy (shamim):

120-105=?

OpenStudy (anonymous):

15

OpenStudy (shamim):

15/120=?

OpenStudy (anonymous):

8 or 0.125?

OpenStudy (shamim):

0.125

OpenStudy (shamim):

Not 8

OpenStudy (anonymous):

so 0.125 is the answer?

OpenStudy (shamim):

No

OpenStudy (shamim):

0.125*100=?

OpenStudy (anonymous):

12.5

OpenStudy (shamim):

12.5% is ur result

OpenStudy (anonymous):

thank you !

OpenStudy (shamim):

Is it clear to u?

OpenStudy (shamim):

Glad to help u!